# m-KAILIN: A Knowledge-Driven Agentic Scientific Corpus Distillation Framework for Biomedical Large Language Models Training

### v2

@@ 4.1.4 Platform Information
Each experiment, either utilizing the API query or performed on an Ubuntu 18.04.6 LTS system featuring an AMD EPYC 7742 processor and four NVIDIA V100 GPUs, was carried out within a Python 3.11.0 environment using PyTorch 2.1.1.
@@ 4.3 Ablation Studies of each Technical Component
Figure 8: Ablation of Two Distinct Question Generation Agents
@@ 4.3.2 Ablation Study on Context Retrieval Agent and Question Evaluation Agent
[…]
@@ 4.5 Case Study of Generated Corpus
2020202030303030404040405050505060606060707070708080808090909090100100100100747474747676767678787878808080808282828284848484Scale (×$\times 10^{3}$)Accuracy (%)m-KAILIN-LLaMA-3 (8B)m-KAILIN-LLaMA-3 (70B)
Figure 12: Evaluation (accuracy(%)) of Llama-3-8B and Llama-3-70B on PubMedQA using datasets of varying scales, from 20,000 (20×$\times 10^{3}$) to 100,000 (100×$\times 10^{3}$).
@@
4.6 Scaling Law of Dataset Distillation
We conducted experiments using datasets of varying scales for continuous pre-training, followed by supervised fine-tuning under consistent experimental settings, to evaluate the impact of different data volumes on model performance. In this study, we selected the Llama-3-8B and Llama-3-70B models as base models, to verify the impact of varying dataset scales on models of different sizes. As shown in Figure 12, we observed that the larger LLaMA-3-70B model demonstrates overall better performance compared to the smaller LLaMA-3-8B model. We attributed this advantage to its larger parameter scale, which provides superior feature capture and generalization capabilities. We also observed that on datasets of varying sizes, both models showed a trend of performance improvement as the dataset size increased. The underlying driver is that larger datasets distilled by m-KAILIN are better able to represent a broader range of biomedical knowledge due to a greater number of documents being comprehensively analyzed. This phenomenon indicates that the m-KAILIN framework indeed produces higher-quality datasets when distilling larger-scale datasets. Moreover, we found that while smaller models may not outperform larger models, they seem to benefit more from progressively larger distilled datasets. We attribute this phenomenon to the degree of alignment between task complexity and model scale. This phenomenon indicates that larger models, due to potential knowledge redundancy, may gain less from incremental datasets on a specific task.

### v3

@@ 4.1.4 Platform Information
Each experiment, either utilizing the API query or performed on an Ubuntu 18.04.6 LTS system featuring an AMD EPYC 7742 processor and eight Cambricon MLU-590B, was carried out within a Python 3.11.0 environment using PyTorch 2.1.1.
((a))
((b))
Figure 7: Evaluations (accuracy (%)) for PubMedQA [jin2019pubmedqa] problems on our models compared to other open-source models, closed-source models, and domain-specific models. Models marked with * indicate domain-specific biomedical models. The # symbol denotes closed-source models.
@@ 4.3 Ablation Studies of each Technical Component
Figure 8: Ablation study on the question generation agent backend. Comparison between using two identical general-purpose agents (Same Agent Backend) and combining a general-purpose agent with a domain-specific agent (Distinct Agent Backend).
@@ 4.3.2 Ablation Study on Context Retrieval Agent and Question Evaluation Agent
4.4 Hyperparameter Experiment
224466886868696970707171727273737474kk (Retrieval Number)Accuracy (%)Accuracy
@@ 4.5 Case Study of Generated Corpus
4.6 Scaling Law of Dataset Distillation
We conducted experiments using datasets of varying scales for continuous pre-training, followed by supervised fine-tuning under consistent experimental settings, to evaluate the impact of different data volumes on model performance. In this study, we selected the Llama-3-8B and Llama-3-70B models as base models, to verify the impact of varying dataset scales on models of different sizes.
20203030404050506060707080809090100100747476767878808082828484Scale ($\times 10^{3}$)Accuracy (%)m-KAILIN-LLaMA-3 (8B)m-KAILIN-LLaMA-3 (70B)
@@ 4.6 Scaling Law of Dataset Distillation
As shown in Figure 12, we observed that the larger LLaMA-3-70B model demonstrates overall better performance compared to the smaller LLaMA-3-8B model. We attributed this advantage to its larger parameter scale, which provides superior feature capture and generalization capabilities. We also observed that on datasets of varying sizes, both models showed a trend of performance improvement as the dataset size increased. The underlying driver is that larger datasets distilled by m-KAILIN are better able to represent a broader range of biomedical knowledge due to a greater number of documents being comprehensively analyzed. This phenomenon indicates that the m-KAILIN framework indeed produces higher-quality datasets when distilling larger-scale datasets. Moreover, we found that while smaller models may not outperform larger models, they seem to benefit more from progressively larger distilled datasets. We attribute this phenomenon to the degree of alignment between task complexity and model scale. This phenomenon indicates that larger models, due to potential knowledge redundancy, may gain less from incremental datasets on a specific task.
9.6%12.2%11.9%11.9%9.6%14.8%15.0%9.2%1989-20002001-20042005-20072008-20092010-20112012-20132014-20152016-2017
@@ 5 Conclusion
[…]
This work is partially supported by the National Natural Science Foundation of China (No. 62506351, No.92470204), the Beijing Municipal Natural Science Foundation (No.4254089). Hengshu Zhu is the corresponding author.